OpenStudy (anonymous):

Need help... Prove: sin (A-B) sin (A+B) = sin^2 A - sin^2 B

OpenStudy (anonymous):

sin(A+B) sin(A -B) = (sinA cosB + cosA sinB) (sinA cosB - cosA sinB) = sin^2A cos^2B - cos^2A sin^2B = sin^2A (1-sin^2B) - cos^2A sin^2B = sin^2A - sin^2B (sin^2A + cos^2A) = sin^2A - sin^2B.

OpenStudy (anonymous):

sin (A-B) sin (A+B) = sin^2 A - sin^2 B LHS=sin (A-B) sin (A+B) =[sinAcosB-cosAsinB][sinAcosB+cosAsinB] \[=\sin^2A \cos^2B-\cos^2A \sin^2B\] =sin ² A ( 1 - sin ² B ) - sin ² B ( 1 - sin ² A ) =sin ² A - sin ² B - sin ² A sin ² B + sin ² A sin ² B =sin ² A - sin ² B = RHS Hence proved

OpenStudy (anonymous):

Thank you!!